Solve each equation by the method of your choice.
2y^2-3y-6=0
and
(x-10)(x-2)=-20

Answer by checkley75(3666)

2y^2-3y-6=0
using the quadratic equation we get:
y=(3+-sqrt[-3^2-4*2*-6])/2*2
y=(3+-sqrt[9+48])/4
y=(3+-sqrt57)/4
y=(3+-7.55)/4
y=(3+7.55)/4
y=10.55/4
y=2.637 answer.
y=(3-7.55)/4
y=-4.55/4
y=1.1375 answer.
----------------------------------------------------------------
(x-10)(x-2)=-20
x^2-12x+20=-20
x^2-12x+20+20=0
x^2-12x+40=0
again using the quadratic equation we get:
x=(12+-sqrt[-12^2-4*1*40])/2*1
x=(12+-sqrt[144-160])/2
x=(12+-sqrt-16)/2
x=(12+-4i)/2
x=12/2+4i/2
x=6+2i answer.
x=(12-4i)/2
x=6-2i answer.